# A HHO formulation for variable density incompressible flows where the density is purely advected

### 04 Mar 2026 v3

@@ 1 Introduction
Several methods for variable density incompressible flows have been proposes in literature. A DG formulation based on exact Riemann solvers for the artificial compressibility perturbation of the problem was proposed by Bassi et al. [3] and assessed in [2]. An entropy–stable DG approximation was devised by Manzanero et al. [22] while a post-processed semi-implicit DG scheme was introduced by [21]. A Discrete Duality Finite Volume (DDFV) scheme was proposed by Goudon et al. [16], while Fu [15] addressed the Cahn-Hilliard phase-field model with a HDG formulation based on divergence-free velocity spaces. Recently, Nordström et al. [23] proposed an energy stable formulation for multi-phase incompressible flows and Dauphin et al. [12] analysed a low-order hybrid method for the variable-density incompressible Navier-Stokes equations. Earlier pioneering works focused on Galerkin discretizations with projection methods, consider for example Guermond et al. [17]-[18].
@@ 3.1 Mesh
The HHO formulation proposed in this work is capable of dealing with two and three dimensional flow problems. Nevertheless, since numerical test cases focus on two space dimensions, discrete settings are provided for the 2D case focusing on simplicial meshes. We consider meshes of the domain $\Omega$ corresponding to couples $\mathcal{M}_{h}\coloneq(\mathcal{T}_{h},\mathcal{F}_{h})$, where $\mathcal{T}_{h}$ is a finite collection of triangular elements such that $h\coloneq\max_{T\in\mathcal{T}_{h}}h_{T}>0$ with $h_{T}$ denoting the diameter of $T$, while $\mathcal{F}_{h}$ is a finite collection of line segments, with $h_{F}$ denoting the length of $F$. It is assumed henceforth that the mesh $\mathcal{M}_{h}$ is shape- and contact-regular, as detailed in [13, Definition 1.4] and that its trace on $\partial\Omega$ is compatible with the partition $\partial\Omega=\partial\Omega_{\rm D}\sqcup\partial\Omega_{\rm N}$. We respectively denote by $\mathcal{F}_{h}^{\rm D}$ and $\mathcal{F}_{h}^{\rm N}$ the sets of Dirichlet and Neumann faces, whereas the set of internal faces is $\mathcal{F}_{h}^{\rm i}\coloneq\mathcal{F}_{h}\setminus\big(\mathcal{F}_{h}^{\rm D}\cup\mathcal{F}_{h}^{\rm N}\big)$. For each mesh cell $T\in\mathcal{T}_{h}$, the faces contained in the cell boundary $\partial T$ are collected in the set $\mathcal{F}_{T}$ and we additionally let $\partial T^{\bullet}\coloneqq\partial T\cap\partial\Omega_{\bullet}$ for $\bullet\in\{{\rm D},{\rm N}\}$. For all $T\in\mathcal{T}_{h}$ and all $F\in\mathcal{F}_{T}$, $\bm{n}_{TF}$ denotes the unit vector normal to $F$ pointing out of $T$ and $\bm{n}_{\partial T}:\partial T\to\mathbb{R}^{d}$ is such that $\bm{n}_{\partial T}|_{F}\coloneqq\bm{n}_{TF}$ for all $F\in\mathcal{F}_{T}$.
@@ 3.4 Semi-discrete HHO formulation
Both in the momentum and the mass conservation equation, advection terms are treated with skew-symmetric formulations see (2)-(3) and (7)-(8), respectively, plus stabilization bilinear forms, see (4) and (9), respectively. We remark that, in the convective term discretization of (2)-(3)-(4), we consider $L^{2}$-projections of degree $k$ for test functions and $L^{2}$-projections of degree $p$, with p=kp{=}k or p=k+1p{=}k{+}1, for the advected velocity unknowns, leading to two different variants of the method. Note that the advective velocity is not projected. In the case p=k+1p{=}k{+}1, since $\bm{u}_{T}\in\mathcal{P}^{k+1}(T)$, it holds $\bm{\pi}_{T}^{p}\bm{u}_{T}=\bm{u}_{T}$, furthermore, since $\bm{u}_{\partial T}\in\mathcal{P}^{k}(\partial T^{{\rm i},{\rm D}})\subset\mathcal{P}^{k+1}(\partial T^{{\rm i},{\rm D}})$ and $\bm{u}_{\partial T}\in\mathcal{P}^{k+1}(\partial T^{{\rm N}})$, it holds $\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}=\bm{u}_{\partial T}$, thus, contrary to the case p=kp{=}k, the formulations is not skew-symmetric. The choice p=kp{=}k, proposed in [4, Remark 3.7], leads to a provably stable formulation, see Sec. 3.7, while the choice p=k+1p{=}k{+}1, proposed in [8], provides improved accuracy in the convection-dominated regime. The diffusion term stabilization of (5), also involving $L^{2}$-projection operators, was first proposed by [20], see also [11]. Note that the HHO discretization of the mass conservation equation in (7) only relies on the cell velocity $\bm{u}_{T}$ which, as result of the HHO discretization of the volume conservation in (10), is $\bm{H}(\operatorname{div})$-conforming and divergence-free, see Sec. 3.5 for details.
@@
$sd^{{\rm slm},\pi^{p}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)$ $\coloneqq sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{p}_{T});\underline{\bm{v}}_{T})-\frac{1}{2}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
@@
We remark that the time derivative in the momentum equation has been modified in view of achieving stability, as first proposed by [17], see Sec. 3.7 for details. We further remark that the second term of (13) is a time derivative stabilization designed to guarantee the robustness of the formulation whenever the normal velocity component vanishes on a mesh face, see Sec. 3.6 for additional details.
@@ 3.5 H(div)-conformity and independence from irrotational body forces
    $\underline{q}_{h}=\left((q_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{q}_{T}\in\mathbb{P}_{d}^{k}(T)$, we get $\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}=0,\;$ i.e., $\nabla\cdot\bm{u}_{T}=0,\forall\bm{x}\in T$, since $\nabla\cdot\bm{u}_{T}\in\mathbb{P}_{d}^{k}(T)$;
@@
3.6 Flux formulation of the mass conservation equation
@@
$\int_{T}z_{T}\bm{u}_{T}\cdot\nabla\rho_{T}=\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{TF}\bigr){\rho}_{T}{z}_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T},$
@@
Notice that, in (18), $\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr){=}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)$, indeed $\rho_{\partial T}$ and ${z}_{\partial T}$ and are single valued over internal faces and $\bm{u}_{T}$ is H-div conforming. Introducing (18) in the definition of $sd^{\rm sms}_{T}$, see (11b), we finally get the following reformulation of the unsteady mass conservation HHO spatial discretization of (13): given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, for all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$,
$sd^{\rm mss}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$ $\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla{z}_{T}+\int_{\partial T}\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})({z}_{T}{-}{z}_{\partial T})$
$+\int_{\partial T^{\rm D}}\left(\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}.$
In the spatial discretization above, the numerical flux reads $\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\coloneqq\pi^{k}_{\partial T}\bigl(\phi^{\rm{td}}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})+\phi^{up}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\bigr)$, with
@@
    $\underline{z}_{h}=\left(({z}_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{z}_{T}\in\mathbb{P}_{d}^{k}(T)$, we obtain the following local mass conservation statement
@@
    |unTF|a, if   unTF0,
    |unTF|b, if   <unTF0.

### 02 May 2026 v4

@@ 1 Introduction
Several methods for variable density incompressible flows have been proposes in literature. A DG formulation based on exact Riemann solvers for the artificial compressibility perturbation of the problem was proposed by Bassi et al. [4] and assessed in [3]. A Volume Of Fluid (VOF) DG formulation for two-phase flows was considered by [24], interestingly, the transport equation for the density is modified introducing a slope-limited $H$-div conforming velocity field serving as advection velocity. An entropy–stable DG approximation was devised by Manzanero et al. [27] while a post-processed semi-implicit DG scheme was introduced by [26]. A Discrete Duality Finite Volume (DDFV) scheme was proposed by Goudon et al. [20], while Fu [19] addressed the Cahn-Hilliard phase-field model with a HDG formulation based on divergence-free velocity spaces. Recently, Nordström et al. [28] proposed an energy stable formulation for multi-phase incompressible flows and Dauphin et al. [16] analysed a low-order hybrid method for the variable-density incompressible Navier-Stokes equations. Earlier pioneering works focused on Galerkin discretizations with projection methods, consider for example Guermond et al. [21]-[22].
@@ 3.1 Mesh
The HHO formulation proposed in this work is capable of dealing with two and three dimensional flow problems. Nevertheless, since numerical test cases focus on two space dimensions, discrete settings are provided for the 2D case focusing on simplicial meshes. We consider meshes of the domain $\Omega$ corresponding to couples $\mathcal{M}_{h}\coloneq(\mathcal{T}_{h},\mathcal{F}_{h})$, where $\mathcal{T}_{h}$ is a finite collection of triangular elements such that $h\coloneq\max_{T\in\mathcal{T}_{h}}h_{T}>0$ with $h_{T}$ denoting the diameter of $T$, while $\mathcal{F}_{h}$ is a finite collection of line segments, with $h_{F}$ denoting the length of $F$. It is assumed henceforth that the mesh $\mathcal{M}_{h}$ is shape- and contact-regular, as detailed in [17, Definition 1.4] and that its trace on $\partial\Omega$ is compatible with the partition $\partial\Omega=\partial\Omega_{\rm D}\sqcup\partial\Omega_{\rm N}$. We respectively denote by $\mathcal{F}_{h}^{\rm D}$ and $\mathcal{F}_{h}^{\rm N}$ the sets of Dirichlet and Neumann faces, whereas the set of internal faces is $\mathcal{F}_{h}^{\rm i}\coloneq\mathcal{F}_{h}\setminus\big(\mathcal{F}_{h}^{\rm D}\cup\mathcal{F}_{h}^{\rm N}\big)$. For each mesh cell $T\in\mathcal{T}_{h}$, the faces contained in the cell boundary $\partial T$ are collected in the set $\mathcal{F}_{T}$ and we additionally let $\partial T^{\bullet}\coloneqq\partial T\cap\partial\Omega_{\bullet}$ for $\bullet\in\{{\rm D},{\rm N}\}$. For all $T\in\mathcal{T}_{h}$ and all $F\in\mathcal{F}_{T}$, $\bm{n}_{TF}$ denotes the unit vector normal to $F$ pointing out of $T$ and $\bm{n}_{\partial T}:\partial T\to\mathbb{R}^{d}$ is such that $\bm{n}_{\partial T}|_{F}\coloneqq\bm{n}_{TF}$ for all $F\in\mathcal{F}_{T}$.
@@ 3.4 Semi-discrete HHO formulation
Both in the momentum and the mass conservation equation, advection terms are treated with skew-symmetric formulations see (2)-(3) and (7)-(8), respectively, plus stabilization bilinear forms, see (4) and (9), respectively. We remark that, in the convective term discretization of (2)-(3)-(4), we consider $L^{2}$-projections of degree $k$ for test functions and $L^{2}$-projections of degree $p$, with p=kp{=}k or p=k+1p{=}k{+}1, for the advected velocity unknowns, leading to two different variants of the method. Note that the advective velocity is not projected. In the case p=k+1p{=}k{+}1, since $\bm{u}_{T}\in\mathcal{P}^{k+1}(T)$, it holds $\bm{\pi}_{T}^{p}\bm{u}_{T}=\bm{u}_{T}$, furthermore, since $\bm{u}_{\partial T}\in\mathcal{P}^{k}(\partial T^{{\rm i},{\rm D}})\subset\mathcal{P}^{k+1}(\partial T^{{\rm i},{\rm D}})$ and $\bm{u}_{\partial T}\in\mathcal{P}^{k+1}(\partial T^{{\rm N}})$, it holds $\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}=\bm{u}_{\partial T}$, thus, contrary to the case p=kp{=}k, the formulations is not skew-symmetric. The choice p=kp{=}k, proposed in [5, Remark 3.7], leads to a provably stable formulation, see Sec. 3.7, while the choice p=k+1p{=}k{+}1, proposed in [10], provides improved accuracy in the convection-dominated regime. The diffusion term stabilization of (5), also involving $L^{2}$-projection operators, was first proposed by [25], see also [15]. Note that the HHO discretization of the mass conservation equation in (7) only relies on the cell velocity $\bm{u}_{T}$ which, as result of the HHO discretization of the volume conservation in (10), is $\bm{H}(\operatorname{div})$-conforming and divergence-free, see Sec. 3.5 for details.
@@
$sd^{{\rm slm},\pi^{p}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)$ $\coloneqq sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{p}_{T});\underline{\bm{v}}_{T})-\frac{1}{2}\int_{\partial T^{{\rm D},{\rm N}}}\rho_{\partial T}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)\bigl(\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
@@
We remark that the time derivative in the momentum equation has been modified in view of achieving stability, as first proposed by [21], see Sec. 3.7 and Sec. 3.6.2 for details. We further remark that the second term of (13) is a time derivative stabilization designed to guarantee the robustness of the formulation whenever the normal velocity component vanishes on a mesh face, see Sec. 3.6.1 for additional details.
@@ 3.5 H(div)-conformity and independence from irrotational body forces
    $\underline{q}_{h}=\left((q_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{q}_{T}\in\mathcal{P}^{k}(T)$, we get $\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}=0,\;$ i.e., $\nabla\cdot\bm{u}_{T}=0,\forall\bm{x}\in T$, since $\nabla\cdot\bm{u}_{T}\in\mathcal{P}^{k}(T)$;
@@
3.6 Flux formulations
3.6.1 Flux formulation of the mass conservation equation
@@
$\int_{T}z_{T}\bm{u}_{T}\cdot\nabla\rho_{T}=\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{TF}\bigr){\rho}_{T}{z}_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T},$ (18)
@@
Introducing (19) in the definition of $sd^{\rm sms}_{T}$, see (11b), we finally get the following reformulation of the global residual of the mass conservation equation defined in (14): for all $\underline{z}_{h}\in\underline{Z}_{h}^{k}$,
$r^{\rm mss}_{h}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h});\underline{z}_{h}\bigr)=\sum_{T\in\mathcal{T}_{h}}\left(\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla{z}_{T}+\int_{\partial T}\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})({z}_{T}{-}{z}_{\partial T})\right)$
${+}\sum_{T\in\mathcal{T}_{h}}\left(\int_{\partial T^{\rm N}}\left(\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}{+}\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}{+}\int_{\partial T^{\rm D}}\left(\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}{+}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}\right)$
$+\sum_{T\in\mathcal{T}_{h}}\left(\frac{1}{2}\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)-\frac{1}{2}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)\right).$
Notice that, since $\rho_{\partial T}$ and ${z}_{\partial T}$ and are single valued over internal faces and $\bm{u}_{T}$ is HH(div)-conforming, it holds
$\sum_{T\in\mathcal{T}_{h}}\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr){=}\sum_{T\in\mathcal{T}_{h}}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr),$ (20)
and, accordingly, the terms on the last line of $r^{\rm mss}_{h}$ cancel. In the spatial discretization residual above, the numerical flux reads $\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\coloneqq\pi^{k}_{\partial T}\bigl(\phi^{\rm{td}}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})+\phi^{up}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\bigr)$, with
@@
    $\underline{z}_{h}=\left(({z}_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{z}_{T}\in\mathcal{P}^{k}(T)$, we obtain the following local mass conservation statement
@@
Note that an internal face is such that $F=\partial T\cap\partial T^{\prime}$, with $^{\prime}\in\mathcal{T}_{h}, TTT\neq T^{\prime}$ and $\bm{n}_{TF}=-\bm{n}_{T^{\prime}F}$. We further remark that $\phi_{\partial T}^{\rm{up}}(\bm{u}_{T}{,}\underline{\rho}_{T})=0$ if the normal velocity component vanishes on a face $F\in\mathcal{F}_{T}$. In this event, thanks to the additional time derivative stabilization provided by $\phi_{\partial T}^{\rm{td}}(\underline{\rho}_{T})$, the face density is still well defined. A similar strategy was proposed to stabilize the linear momentum equations in the limit of vanishing viscosity, see [5]. Note that, disregarding the time derivative stabilisation, i.e.,  $\phi_{\partial T}^{\rm{td}}(\underline{\rho}_{T})=0$, $\forall F\in\mathcal{F}_{h}^{\rm i}$, we simply get